# A Survey on Failure Analysis and Fault Injection in AI Systems

### v1

@@ 9.1.2. Network
Network failures have long been a significant area of research, particularly concerning traditional faults such as congestion, packet loss, and latency, which have been extensively discussed and studied. Hassan et al. (Khan et al., 2023) study how network faults occurring in the links between the nodes of the cloud management platforms can propagate and affect the applications that are hosted on the virtual machines.
However, with the advancement of AI systems, the networking infrastructure of AI systems has become increasingly complex, leading to the emergence of unique fault types. Distributed deep learning training across multiple compute nodes is pretty common and these nodes are internally interconnected with a high-speed network (e.g., via InfiniBand). Gao et al. (Gao et al., 2023b) and C4 (Dong et al., 2024) classify network faults on AI platform into InfiniBand-related and Ethernet-related.
@@ 9.1.3. Node
A node in AI platform is a distinct schedulable unit for computation with GPUs, CPUs, main memory, disks, and network. Gao et al. (Gao et al., 2023b) classify node faults on AI platform into node outage, node damage and node preemption. These faults summarize the impact caused by faults occurring inside the node. In addition to these faults, communication faults between nodes are also of concern (Chen et al., 2020a). Thus, we classify node faults into two types, namely node crash, node partition.
@@
Tool Description Instr. Link
@@ 9.2.1. Hardware Accelerator
  • Software. At present, various FI techniques exist at different levels of programming languages. Commonly, faults are injected at the GPU assembly code (SASS) level, which is the instruction-level code running directly on the GPU. For instance, SASSIFI (Hari et al., 2017) employs the SASSI (SASS Instrumentation) framework for compile-time instrumentation of SASS code to insert fault injection code. GPU-Qin (Fang et al., 2014) utilizes CUDA-GDB to control faults during runtime without modifying the code. NVBitFI (Tsai et al., 2021) dynamically loads relevant code as a library during runtime for fault injection. Besides SASS level, there are also works at the PTX and LLVM IR levels. For instance, LLFI-GPU (Li et al., 2016) improves FI in LLVM IR, the intermediate representation language.
  • Hardware/Simulation. Hardware or simulation-based approaches provide a more realistic reflection of fault impacts. Direct methods include radiation experiments to evaluate hardware reliability. For example, Oliveira et al. (Oliveira et al., 2017) use beam tests to study the radiation effects on NVIDIA and Intel accelerators, quantifying and limiting radiation effects by observing amplitude and fault propagation in final outputs. Simulation approaches replace actual hardware faults (e.g., electromagnetic interferences at the physical level) by injecting their expected effects on memory and registers (e.g., flipped and stuck bits), thus approximating the hardware fault process (Zhang et al., 2019a; Deng et al., 2015b; Temam, 2012). They simulate faults at different levels, such as modifying simulator variables, introducing faults at the RTL level, and injecting faults at the gate level. The NVIDIA Data Center GPU Manager (DCGM) includes an fault injection framework allows users to simulate the fault handling behavior of the DCGM APIs when GPU faults are encountered (NVIDIA, 2024b).
@@
  • Reconfiguration-based techniques. In reconfiguration-based techniques, faults are injected by changing the bit stream needed for configuring FPGA. Antoni et al. introduce a novel methodology for injecting single event upsets (SEUs) in memory elements. This approach involves conducting the injection directly within the reconfigurable FPGA, leveraging the runtime reconfiguration capabilities of the device (Antoni et al., 2002). Gabriel et al. propose a fault injection tool to evaluate the impact of faults in an FPGA’s configuration memory (Nazar and Carro, 2012).
@@
TPU. TPU, as a variant of systolic arrays, represents a parallel computing architecture. The intrinsic parallelism and matrix multiplication efficiency inherent in systolic arrays empower them to achieve superior performance in both the training and inference phases of deep neural networks. Numerous studies have been conducted to investigate faults associated with systolic arrays. Udit et al. (Agarwal et al., 2023) propose an RTL-level fault injection framework for systolic arrays. Using this framework, they characterized the software effect of faults induced by stuck-at faults within the multiply and accumulation units of the systolic array. Zhang et al. (Zhang et al., 2020a) and Holst et al. (Holst et al., 2021) study the effects of timing faults in systolic arrays, thus, degrading DNN’s accuracy.
@@ 9.2.2. Network
In addition to general purpose fault injection tools such as ChaosBlade (Chaosblade, 2024) that can introduce common network faults such as network delay and network packet loss, there are now some network fault injection tools for large infrastructure. Domenico et al. propose ThorFI (Cotroneo et al., 2022), a novel approach for virtual networks in infrastructures. ThorFI is designed to provide non-intrusive fault injection capabilities for a cloud tenant, and to isolate injections from interfering with other tenants on the infrastructure. Michal et al. propose NetLoiter (Rozsíval and Smrcka, 2023), which can introduce real world fault, including lossy channels, network jitter, data corruption, or disconnections.
@@ 9.3. Gap between Failure Analysis and Fault Injection in AI Infrastructure
From the Table 18, several key observations regarding the gap between FA and FI at the infrastructure level in AI systems can be made:
Limited coverage of fault types. The table shows that several types of faults, such as ”Out of Memory”, ”Off the Bus”, ”InfiniBand Fault”, and ”Ethernet Fault”, are not currently being simulated by any of the listed FI tools. This suggests a significant gap in the ability of current tools to simulate a comprehensive set of failure scenarios at the infrastructure level in AI systems.
Hardware and network-specific FI The existing tools seem to focus on specific areas of the infrastructure. For example, ”Bit-flip Fault” and ”Stuck-at Fault” are well-covered by tools designed for hardware faults like SASSIFI, LLFI-GPU, and ThunderVolt. On the other hand, network-related faults such as ”Network Jam”, ”Network Loss” are covered by NetLoiter, FCatch. This suggests that the current fault injection tools are specialized for either hardware or network faults but not both.
Lack of realism in hardware accelerator faults. Most of the existing fault injection tools for hardware accelerators, such as GPUs, operate at the software level or are based on simulations. The faults generated by these methods can be classified as emulated faults. While emulation provides high efficiency, a significant drawback is that the faults may lack realism. This is because emulated faults may not accurately represent the complex physical processes that cause real hardware faults. As a result, the conclusions drawn from studies using these tools may not fully apply to real-world scenarios where hardware faults occur. This lack of realism in emulated faults represents another significant gap in the current state of fault injection for AI systems.
@@
Failure SASSIFI LLFI-GPU SCFIT NVBitFI ThunderVolt ThorFI NetLoiter FCatch CoFI Covered
Bit-flip Fault True
@@
Node Crash True
@@ 10. Future Opportunities of Fault Injection in AI Systems
FI is a widely used technique for evaluating the reliability of AI systems. However, as discussed above, there is a huge gap between FA and FI. Bridging this gap is a research opportunity in future.
10.1. More Comprehensive Fault Injection
Support for more faults types. As shown in previous sections, there are a number of real faults that have occurred that cannot yet be simulated by existing FI tools. For example, no FI tool in Table 14 can simulate GPU contention faults in Table 13. Ignoring these faults leads to incomplete FI testing and may hide risks in AI systems. Therefore, based on the FA results in this paper, designing FI tools that cover as much as possible all the faults that have occurred historically can provide a more comprehensive assessment of the reliability and fault tolerance of AI systems.
Cross-layer multiple fault injection. Current FI tools typically only inject a fault at a single layer. However, in distributed systems, there are situations where multiple faults occur simultaneously (Yu et al., 2021). Designing tools that support simultaneous injection of multiple faults can facilitate the creation of more complex fault scenarios. At the same time, considering the dependencies between AI system layers can also enable the simulation of richer fault scenarios through cross-layer fault linkage injection. Consequently, the development of tools that facilitate cross-layer multiple fault injection will prove advantageous in evaluating the reliability and fault-tolerance of AI systems when multiple faults occur at different layers simultaneously.
LLM-specific FI tool. As LLMs continue to gain prominence in both academic and industrial sectors, the importance of assessing their reliability cannot be overstated. Given the unique characteristics and potential failure of LLMs, there is a pressing need for the development of FI tools specifically designed for these systems. Such tools should be capable of simulating LLM-specific failures, including those related to language understanding, reasoning, and generation.
10.2. More Generalizable Fault Injection
Compatible with more layers and frameworks. From the perspective of FI generality, current FI tools are in a state of fragmentation. For example, PytorchFI (Pytorchfi, 2024) can only inject faults related to PyTorch, while TensorFI (DependableSystemsLab, 2024a) can only inject faults to TensorFlow. Even for the same AI framework, there may be conceptual differences between versions. For example, TensorFlow 1 and TensorFlow 2 exhibit significant differences in API usage and runtime logic, requiring separate fault injection tools (e.g., TensorFI (DependableSystemsLab, 2024a) and InjectTF (Mbsa-tud, 2024)) to be designed for them. This results in a considerable number of FI tools that engineers must maintain, as well as a significant amount of time required to learn to use them. Consequently, the design of a more unified tool that can inject faults across different layers and across different frameworks is of great importance in the future.
Non-instrumented injection. Numerous contemporary FI tools in Table 8 and Table 11 necessitate the instrumentation of the target for FI (e.g., modifying the framework source code). This increases the difficulty of utilising FI tools. Given that the majority of frameworks and algorithms associated with AI systems are implemented in Python, it is possible to implement python bytecode modifications that do not necessitate code instrumentation, as was previously the cases (Kmaork, 2024; Yu et al., 2023a). Even for non-Python implementations, it is possible to achieve non-instrumented FI through eBPF (Yu et al., 2023b; Trailofbits, 2024). Future research into work injection without code instrumentation could facilitate the development of more user-friendly FI tools.
10.3. More Intelligent Fault Injection
@@
A FI policy must specify the location of the FI, the type and intensity of the fault, and so forth. This combination of several attributes forms a vast search space for FI policies. The objective is to identify valuable FI policies from this vast search space and to discover as many faults as possible with as few fault injections as possible. Currently, this process relies mainly on expert experience, resulting in high and inefficient labour costs. In the future, it is anticipated that intelligent algorithms will be introduced to facilitate the selection of fault injection strategies in an intelligent manner.
LLM-based Fault Injection. LLM has already demonstrated its capabilities in several software engineering tasks (Pearce et al., 2023; Wei et al., 2023; Xia et al., 2023). In the future, the integration of LLM and Reinforcement Learning from Human Feedback (RLHF) will enable the translation of natural language descriptions of fault scenarios directly into FI policies, thereby reducing the manual effort required to design and implement fault scenarios in AI systems (Cotroneo and Liguori, 2024).
@@ 11. Conclusion
In this study, we have examined the current state of FA and FI in AI systems, providing a critical overview of prevalent failures, the capabilities of existing FI tools, and the gaps between simulated and actual failures. Our analysis, based on a thorough review of relevant paper and code repositories, has revealed significant gaps in the ability of current FI tools to simulate the wide range of failures that occur in real-world AI systems. Moreover, this survey contributes by discussing technical challenges of FI in AI systems and outlining future research avenues. The findings of this study serve as a foundation for further advancements in the field of FA and FI for AI systems.

### v2

@@ 9.1.2. Network
Network failures have long been a significant area of research, particularly in relation to traditional faults such as congestion, packet loss, and latency, which have been extensively discussed and studied. Hassan et al. (Khan et al., 2023) study how network faults occurring in the links between the nodes of the cloud management platforms can propagate and affect the applications that are hosted on the virtual machines.
However, with the advancement of AI systems, the networking infrastructure of AI systems has become increasingly complex, leading to the emergence of unique fault types. Distributed deep learning training across multiple compute nodes is pretty common, and these nodes are internally interconnected with a high-speed network (e.g., via InfiniBand). Gao et al. (Gao et al., 2023b) and C4 (Dong et al., 2024) classify network faults on the AI platform into InfiniBand-related and Ethernet-related.
@@ 9.1.3. Node
A node on the AI platform is a distinct schedulable unit for computation with GPUs, CPUs, main memory, disks, and networks. Gao et al. (Gao et al., 2023b) classify node faults on AI platform into node outage, node damage, and node preemption. These faults summarize the impact caused by faults occurring within the node. In addition to these faults, communication faults between nodes are also of concern (Chen et al., 2020a). Thus, we classify node faults into two types, namely node crash and node partition.
@@
Tool Description Instrument Link
@@ 9.2.1. Hardware Accelerator
  • Software. At present, various FI techniques exist at different levels of programming languages. Typically, faults are injected at the GPU assembly code (SASS) level, which is the instruction-level code running directly on the GPU. For instance, SASSIFI (Hari et al., 2017) employs the SASSI (SASS Instrumentation) framework for compile-time instrumentation of SASS code to insert fault injection code. GPU-Qin (Fang et al., 2014) utilizes CUDA-GDB to control faults during runtime without modifying the code. NVBitFI (Tsai et al., 2021) dynamically loads relevant code as a library during runtime for fault injection. Besides SASS level, there are also works at the PTX and LLVM IR levels. For instance, LLFI-GPU (Li et al., 2016) improves FI in LLVM IR, the intermediate representation language.
  • Hardware/Simulation. Hardware or simulation-based approaches provide a more realistic reflection of fault impacts. Direct methods include radiation experiments to evaluate hardware reliability. For example, Oliveira et al. (Oliveira et al., 2017) use beam tests to study the radiation effects on NVIDIA and Intel accelerators, quantifying and limiting radiation effects by observing amplitude and fault propagation in final outputs. Simulation approaches replace actual hardware faults (e.g., electromagnetic interferences at the physical level) by injecting their expected effects on memory and registers (e.g., flipped and stuck bits), thus approximating the hardware fault process (Zhang et al., 2019a; Deng et al., 2015; Temam, 2012). They simulate faults at different levels, such as modifying simulator variables, introducing faults at the RTL level, and injecting faults at the gate level. The NVIDIA Data Center GPU Manager (DCGM) includes an fault injection framework allows users to simulate the fault handling behavior of the DCGM APIs when GPU faults are encountered (NVIDIA, 2024b).
@@
  • Reconfiguration-based Techniques. In reconfiguration-based techniques, faults are injected by changing the bit stream needed to configure the FPGA. Antoni et al. introduce a novel methodology for injecting single event upsets (SEUs) in memory elements. This approach involves conducting the injection directly within the reconfigurable FPGA, leveraging the runtime reconfiguration capabilities of the device (Antoni et al., 2002). Gabriel et al. propose a fault injection tool to evaluate the impact of faults in an FPGA’s configuration memory (Nazar and Carro, 2012).
@@
TPU. TPU, as a variant of systolic arrays, represents a parallel computing architecture. The intrinsic parallelism and matrix multiplication efficiency inherent in systolic arrays empower them to achieve superior performance in both the training and inference phases of deep neural networks. Numerous studies have been conducted to investigate faults associated with systolic arrays. Udit et al. (Agarwal et al., 2023) propose an RTL-level fault injection framework for systolic arrays. Using this framework, they characterized the software effect of faults induced by stuck faults within the multiply and accumulation units of the systolic array. Zhang et al. (Zhang et al., 2020a) and Holst et al. (Holst et al., 2021) study the effects of timing faults in systolic arrays, thus degrading DNN accuracy.
@@ 9.2.2. Network
In addition to general purpose fault injection tools such as ChaosBlade (Chaosblade, 2024) that can introduce common network faults such as network delay and network packet loss, there are now some network fault injection tools for large infrastructure. Domenico et al. propose ThorFI (Cotroneo et al., 2022), a novel approach for virtual networks in infrastructures. ThorFI is designed to provide non-intrusive fault injection capabilities for a cloud tenant, and to isolate injections from interfering with other tenants on the infrastructure. Michal et al. propose NetLoiter (Rozsíval and Smrcka, 2023), which can introduce real-world faults, including lossy channels, network jitter, data corruption, or disconnections.
@@ 9.3. Gap between Failure Analysis and Fault Injection in AI Infrastructure
From the Table 19, several key observations can be made about the gap between FA and FI at the infrastructure level in AI systems.
Limited Coverage of Fault Types. The table shows that several types of faults, such as “Out of Memory”, “Off the Bus”, “InfiniBand Fault” and “Ethernet Fault”, are not currently being simulated by any of the listed FI tools. This suggests a significant gap in the ability of current tools to simulate a comprehensive set of failure scenarios at the infrastructure level in AI systems.
Hardware and Network-specific FI The existing tools seem to focus on specific areas of the infrastructure. For example, “Bit-flip Fault” and “Stuck-at Fault” are well-covered by tools designed for hardware faults such as SASSIFI, LLFI-GPU, and ThunderVolt. On the other hand, network-related faults such as “Network Jam”, “Network Loss” are covered by NetLoiter, FCatch. This suggests that current FI tools are specialized for hardware or network faults but not both.
Lack of Realism in Hardware Accelerator Faults. Most existing FI tools for hardware accelerators, such as GPUs, operate at the software level or are based on simulations. The faults generated by these methods can be classified as emulated faults. Although emulation provides high efficiency, a significant drawback is that the faults may lack realism. This is because emulated faults may not accurately represent the complex physical processes that cause real hardware faults. As a result, the conclusions drawn from studies using these tools may not fully apply to real-world scenarios where hardware faults occur. This lack of realism in emulated faults represents another significant gap in the current state of fault injection for AI systems.
@@
Group Failure SASSIFI LLFI-GPU SCFIT NVBitFI ThunderVolt ThorFI NetLoiter FCatch CoFI Covered
Hardware Accelerator Bit-flip Fault True
@@
Node Node Crash True
@@
10. Threats to Validity
To ensure the rigor and credibility of our systematic survey, we acknowledge and address potential threats to validity. These threats are categorized into study selection validity and data validity, and we outline the mitigation strategies employed to minimize their impact.
10.1. Study Selection Validity
One inherent threat is the possibility of not including all relevant studies in the field. To mitigate this, we adopted a multi-phase approach:
  1. (1)
    Search String Development. We conducted multiple pilot tests to refine the search string, ensuring it captured a comprehensive set of seed studies. Four researchers independently performed preliminary searches, and the final search string was selected through a consensus meeting. The final search string is shown in Section 3.1.
  2. (2)
    Bidirectional Snowballing. To address sampling bias and publication bias, we employed iterative backward and forward snowballing, as described in Section 3.1.2. This approach allowed us to identify 142 primary studies, ensuring a broad coverage of failure analysis (FA) and fault injection (FI) research in AI systems. Snowballing also mitigated the threat of missing relevant studies due to inconsistent terminology or gaps in the search string, which is a common limitation of string-based search methodologies.
  3. (3)
    Inclusion/Exclusion Criteria. To minimize bias in the selection of primary studies, we defined and iteratively refined the inclusion/exclusion criteria in our protocol. Each study was independently evaluated by two researchers from different institutions. Disagreements were resolved through discussion and a third researcher was involved if consensus could not be reached. This rigorous process ensured objectivity and reproducibility.
10.2. Data Validity
Threats related to data extraction and classification were addressed as follows.
  1. (1)
    Data Extraction Protocol. Two researchers independently extracted data from each primary study using a predefined data extraction form. Regular consensus meetings were held to resolve discrepancies and refine the extraction criteria.
  2. (2)
    Classification Process. For classifying studies into the six AI layers (Service, Model, Framework, Toolkit, Platform, Infrastructure), we followed an iterative process:
    • The initial classifications were performed independently by two researchers.
    • Regular meetings were held to discuss and resolve ambiguous cases.
    • A third researcher was consulted for unresolved disagreements.
    The inter-rater reliability for classification was measured using Cohen’s Kappa (k = 0.82), indicating high consistency.
  3. (3)
    Domain Expert Validation. To further enhance the validity of our findings, we consulted three domain experts with over 10 years of experience in AI systems and software engineering. Their feedback was incorporated into the final classification and analysis.
@@ 11. Future Opportunities of Fault Injection in AI Systems
Fault injection is a widely used technique for evaluating the reliability of AI systems. However, as discussed in previous sections, there are significant gaps between FA and FI. Bridging these gaps presents critical research opportunities to advance FI techniques in AI systems. In the following, we outline concrete and reliable directions for future work, prioritizing the resolution of identified shortcomings to enhance the efficacy of fault injection.
11.1. Bridging the Gap: Enhanced Fault Injection Techniques
This section outlines specific techniques to address the limitations of current FI methods, aligning them more effectively with the findings of failure analysis.
Expanding Fault Type Coverage with Targeted Injection Strategies. Based on the gaps between FA and FI presented earlier, there are several fault types that are not adequately covered by existing FI tools. To address this, future research should focus on developing targeted injection strategies for these specific fault types.
  • GPU Contention Faults: With the rising prominence of GPU-based AI systems, it is essential to devise FI techniques that simulate GPU resource contention. Techniques could involve the introduction of artificial delays or constraints on GPU memory and computational resources during the training or inference phases of models.
  • Memory Leaks: Developing fault injection modules that intentionally induce controlled memory leaks within AI frameworks or toolkits will facilitate evaluation of the system’s capability to detect and recover from memory-related malfunctions.
  • Configuration Faults: Designing FI mechanisms to deliberately introduce misconfigurations in AI platform settings, such as incorrect parameter specifications or incompatible library versions, will assist in evaluating system robustness against environmental inconsistencies.
Cross-Layer and Multi-Fault Injection with Dependency Awareness. Contemporary FI tools typically inject faults in isolation, while real-world failures often stem from multiple interrelated faults across different layers. Future solutions should encompass.
  • Simultaneous Injection of Multiple Faults: Current FI tools typically only inject a fault in a single layer. However, in distributed systems, there are situations in which multiple faults occur simultaneously (Yu et al., 2021). Enabling concurrent injection of multiple faults will facilitate simulation of complex failure scenarios inherent in AI systems. This approach reflects the multifaceted nature of real-world faults, improving predictive reliability.
  • Cross-Layer Fault Linkage Injection: Considering the dependencies between AI system layers can also enable the simulation of richer fault scenarios through cross-layer fault linkage injection. Future FI tools should be equipped to consider layer dependencies, allowing simulations of cascading failures. For example, a fault in the data preprocessing layer may lead to subsequent failures in model training. Such cross-layer fault linkage enables more comprehensive evaluations of system reliability and fault tolerance.
11.2. Future Directions: Intelligent and Generalizable Fault Injection
This section explores future development directions for FI, focusing on intelligent automation and increased generality.
@@ 11.2. Future Directions: Intelligent and Generalizable Fault Injection
Intelligent Fault Injection Policy Generation. A foundational FI policy must encapsulate numerous attributes, including fault location, type, and intensity. This combination of several attributes forms a vast search space for FI policies. The objective is to identify valuable FI policies from this vast search space and to discover as many faults as possible with as few fault injections as possible. Currently, this process relies mainly on expert experience, resulting in high and inefficient labor costs. In the future, intelligent algorithms are anticipated to be introduced to facilitate the selection of fault injection strategies in an intelligent manner. LLM has already demonstrated its capabilities in several software engineering tasks (Pearce et al., 2023; Wei et al., 2023; Xia et al., 2023). In the future, the integration of LLM and human feedback reinforcement learning (RLHF) will enable the translation of natural language descriptions of fault scenarios directly into FI policies, thus reducing the manual effort required to design and implement fault scenarios in AI systems (Cotroneo and Liguori, 2024).
Compatible with More layers and Frameworks. From the perspective of FI generality, the current FI tools are in a state of fragmentation. For example, PytorchFI (Pytorchfi, 2024) can only inject faults related to PyTorch, while TensorFI (DependableSystemsLab, 2024a) can only inject faults to TensorFlow. Even for the same AI framework, there may be conceptual differences between versions. For example, TensorFlow 1 and TensorFlow 2 exhibit significant differences in API usage and runtime logic, requiring separate fault injection tools (e.g., TensorFI (DependableSystemsLab, 2024a) and InjectTF (Mbsa-tud, 2024)) to be designed for them. This results in a considerable number of FI tools that engineers must maintain, as well as a significant amount of time required to learn to use them. Consequently, the design of a more unified tool that can inject faults across different layers and across different frameworks is of great importance in the future.
Non-instrumented Injection. Numerous contemporary FI tools in Table 9 and Table 12 necessitate the instrumentation of the target for FI (e.g., modifying the framework source code). This increases the difficulty of utilising FI tools. Given that the majority of frameworks and algorithms associated with AI systems are implemented in Python, it is possible to implement Python bytecode modifications that do not necessitate code instrumentation, as previously the cases (Kmaork, 2024; Yu et al., 2023a). Even for non-Python implementations, it is possible to achieve non-instrumented FI through eBPF (Yu et al., 2023b; Trailofbits, 2024). Future research into work injection without code instrumentation could facilitate the development of more user-friendly FI tools.
@@ 12. Conclusion
In this study, we have examined the current state of FA and FI in AI systems, providing a critical overview of prevalent failures, the capabilities of existing FI tools, and the gaps between simulated and actual failures. Our analysis, based on a thorough review of 142 studies, has revealed significant gaps in the ability of current FI tools to simulate the wide range of failures that occur in real-world AI systems. Moreover, this survey contributes by discussing technical challenges of FI in AI systems and outlining future research avenues. The findings of this study serve as a foundation for further advancements in the field of FA and FI for AI systems.
Acknowledgements.
We greatly appreciate the insightful feedback from the anonymous reviewers. This work was supported in part by the National Natural Science Foundation of China under Grant 62272495 and the Guangdong Basic and Applied Basic Research Foundation under Grant 2023B1515020054. The corresponding author is Pengfei Chen.